Suppose that an emergency medicine unit registers on the average 3 patients with paracot poisoning per week. Calculate the probability that in a given week the unit will register, (i) one or more cases of paracot poisoning, (ii) two or more cases but less than 5 cases of paracot poisoning


1
Expert's answer
2022-01-07T14:25:59-0500

Let "X="the number of patients with paracot poisoning registered per week: "X\sim Po(\lambda t)."

Given "\lambda t=3"

(i)


"P(X\geq 1)=1-P(X=0)=1-\dfrac{e^{-3}(3)^0}{0!}"

"=1-e^{-3}\approx0.950213"

(ii)


"P(2\leq X<5)=P(X=2)+P(X=3)"

"+P(X=4)=\dfrac{e^{-3}(3)^2}{2!}+\dfrac{e^{-3}(3)^3}{3!}+\dfrac{e^{-3}(3)^4}{4!}"

"=e^{-3}(\dfrac{9}{2}+\dfrac{9}{2}+\dfrac{27}{8})=\dfrac{99e^{-3}}{8}\approx0.616115"